# thanks: Corresponding author

### v1

@@ IV Dependence of partial SFF on $\tilde{M}$
To complete our analysis of the pSFF, we investigate the fluctuation in pSFF due to the choice of subsystem. We compute pSFF for $L_{A}=2$ subsystems over $200$ graph realizations for multiple $L$. In each realization, we choose all possible subsystems consisting of two sites as $A$ and the rest of the sites constitute $B$. Figure 6(a) shows the distribution of asymptotic pSFF values for $L_{A}=2$ and multiple $L$ at $\tilde{M}=0.6$. The variance reduces with increasing $L$. At large $L$, the choice of sites does not affect the asymptotic pSFF in the chaotic regime. The peak also shows a shift towards the expected value (in chaotic spectra) $\approx 1/d_{A}^{2}$ ($0.0625$ for $L_{A}=2$) with increasing $L$. This behavior contrasts sharply with the near-integrable regime ($\tilde{M}=0.9$), where ETH breaks down. Here, Fig. 6(b) shows a broad, multi-peaked distribution for small $L$. The distribution gets sharper with $L$ for fixed $L_{A}$. However, is is significantly broader than the pSFF distribution in the chaotic case for the same $L,L_{A}$.
@@
As the pSFF can be experimentally obtained using randomized measurement, we therefore briefly comment on the number of measurements required to estimate pSFF with sufficient accuracy to capture the physics described in this section. As established in Ref [40], the relative error, $\epsilon$, of the pSFF measurement scales with the number of experimental shots, $N_{m}$, as $\epsilon\propto\sigma_{A}/{K_{A}\sqrt{N_{m}}}$ at each instant of time where $\sigma_{A}$ is the variance in the shadow estimator of pSFF. This implies that the number of measurements needed to reach a desired precision, $N_{m}\propto\sigma_{A}^{2}/K_{A}^{2}$. The number of required measurement scales inversely with the instantaneous value of pSFF. This makes the estimation of the dip in pSFF (i.e around $t=t_{Th}$) challenging for large subsystem in chaotic regime as $K_{A}(t\approx t_{Th})=O(10^{-3})$ in our model. Compared to this, the asymptotic value of pSFF serves as a better distinguishing factor as the magnitude is typically one order of magnitude larger here ($t>t_{p}$). However, obtaining plateau requires long time evolution, making it susceptible to decoherence. The correlation hole appears at an earlier time, though it requires a larger number of measurements.
@@ V Operator delocalization through Krylov complexity
Unitary evolution of a closed quantum system with a generic Hamiltonian results in initial local operators growing into more complex operators with time. While probes like Out-of-Time-Ordered Correlators (OTOCs) [21] can be studied for operator spreading, their reliance on spatial separation is ill-suited for random graphs that lack an inherent local structure. A locality-independent quantification of complexity is the measure of the extent to which an operator spreads in a basis composed of increasingly complicated operators under time evolution [64, 65, 66]. In this section, we analyze the dynamics of operator spreading in the system through Krylov complexity. The time evolution of an operator is generated via application of the propagator constructed with the Liouvillian superoperator, and the resulting KC dynamics provide a diagnostic to quantify chaos and localization.
@@
where Liouvillian, $\mathcal{L}=(\mathbb{I}\otimes H-H^{T}\otimes\mathbb{I})$ and $|\mathcal{O})=\mathrm{vec}(\mathcal{O})\in\mathcal{H}_{d}\otimes\mathcal{H}_{d}$ refers to the vectorized operator defined using,
@@
The dimension of $\mathcal{K}_{\mathcal{O}}$ is denoted as $K$ and bounded [68] as $1\leq K\leq d^{2}-d+1$. Evidently, the nested commutator operators in $\mathcal{K}_{\mathcal{O}}$ gets more wide (have larger support) with increasing order. We define an inner product on operator space as $(A|B)=\mathrm{Tr}(A^{\dagger}B)$, and work with a complete orthonormal basis $\{|\mathcal{O}_{n})\}_{n=0}^{K-1}$ in $\mathcal{K}_{O}$ obtained using and Lanczos algorithm implemented through succesive full Gram-Schmidt orthogonalization (See Appendix A). Here, $|\mathcal{O}_{0})=\mathrm{vec}(\mathcal{O}(0))$ i.e. the first element of the basis is the initial operator. This orthonormal basis in operator space is denoted as Krylov basis. The Liouvillian takes a tri-diagonal form when expanded in $\{|\mathcal{O}_{n})\}$,
@@
where diagonal elements are 0 and off-diagonal elements ($\{b_{n}\}$}) form an ordered set of coefficients referred to as Lanczos coefficients. Under time-evolution with $\mathcal{L}$,
@@
where $n\in[0,K-1]$ and $\phi_{n}(0)=\delta_{n,0}$ is taken as initial condition, $b_{0}=b_{K}=0$ is the boundary condition. The form of $\mathcal{L}$ suggests an interpretation of this dynamics analogous to a particle hopping problem on a semi-infinite chain, with $\{|O_{n})\}_{n=0}^{K-1}$ acting as the sites and the hopping amplitudes given by the Lanczos coefficients. To elucidate this, we can represent $\mathcal{L}$ in Krylov basis,
$\mathcal{L}=\sum_{n=0}^{K-1}b_{n}(|\mathcal{O}_{n})(\mathcal{O}_{n-1}|+|\mathcal{O}_{n-1})(\mathcal{O}_{n}|)$ (14)
@@
Figure 7: Analysis of the Lanczos coefficients, $b_{n}$, for an initial operator with support $|\mathcal{O}|=1$. The main panel shows the average coefficients, $\overline{b_{n}}$, as a function of the iteration number $n$ for $\tilde{M}=0.6,0.9$. Error bars represent the standard error of the mean. ( Inset )(a) shows the initial growth of $\overline{b_{n}}$ for the first 20 iterations. (b) shows the distribution of the logarithm of the ratio of consecutive coefficients, $\log(\overline{x})$, where $x_{n}=b_{n}/b_{n+1}$. (c) shows showing the variance of the $\log(\overline{x})$ distribution as a function of $\tilde{M}$. All data are for a system of size $L=6$ with Hamiltonian parameters $g=1.0$ and $h=0.1$, averaged over $500$ graph realizations.
@@
The corresponding evolution of wavefunction representing this particle, $\mathbb{\phi}=\{\ \phi_{0},\imath\phi_{1},\cdots,\imath^{K-1}\phi_{K-1}\}$ can be obtained by solving the set of coupled equations in 13. The KC is defined as,
$\mathcal{C}(t)=\sum_{n=0}^{K-1}n|\phi_{n}(t)|^{2}$ (15)
@@
measuring the particle’s mean position in the Krylov chain. Alternatively, $|\mathcal{O}_{n}\rangle$ involves $n$-th order nested commutators, resulting in Krylov basis elements that have increasingly larger support and more complicated structure with increasing $n$. KC captures the spread of a time-evolved operator in this Krylov basis. We denote the support of an operator $\mathcal{O}$ as $|\mathcal{O}|$.
@@
The Lanczos coefficients $b_{n}$ control how quickly operators spread through the Krylov basis during time evolution. The scaling of $b_{n}$ with $n$ and its fluctuation serve as a signature that can distinguish quantum chaotic dynamics from integrable or localized dynamics [69]. In the following discussion, we take the initial operator in a simple product form.
$\mathcal{O}(0)=\prod_{i=1}^{|\mathcal{O}|}Y_{i}$ (16)
Here, $|\mathcal{O}|$ denotes the size of $\mathcal{O}$ i.e. number of sites where the operator acts non-trivially. Note that we do not have any underlying spatial structure in the lattice; thus, no obvious local structure can be attributed to $\mathcal{O}(0)$ when it has support larger than $1$. In Fig. 7, we show the distribution of $\overline{b_{n}}$ as a function of $n$. $\overline{b_{n}}$ grows at small $n$, saturates at intermediate $n$ and starts descending finally. In a single particle picture, this suggests a hopping problem with site-dependent hopping parameters. The fluctuation in Lanczos amplitude grows as $\tilde{M}\to 1$ i.e., close to the integrable limit. To quantify this fluctuation, we probe the distribution of the ratio of consecutive Lanczos coefficients, $x_{n}=\left(\frac{b_{n}}{b_{n+1}}\right)$. The distribution of $x$ has a mean 0 with the variance increasing as $\tilde{M}\to 1$. In the single particle picture, this indicates the fluctuation in the hopping parameter for adjacent bonds goes up as we approach integrability [70] at extreme $\tilde{M}$. The inset (a) in Fig. 7 shows the growth and subsequent saturation of Lanczos coefficients for small $n$. In the chaotic regime, this saturation is caused due to finite system sizes.
@@
Figure 8: First few Lanczos coefficients and Early-time growth of Krylov complexity. (a) Log-log plot of the average Lanczos coefficients, $\overline{b_{n}}$, versus the iteration number $n$ for several system sizes $L$ at $\tilde{M}=0.6$. Dashed lines show power-law fits, $\overline{b_{n}}\propto n^{\delta}$. All data points are averaged over $200$ graph realizations. (b) The corresponding early-time growth of Krylov complexity, $\overline{\mathcal{C}}(t)$, for the chaotic ($\tilde{M}=0.4$, red) and near-integrable ($\tilde{M}=0.9$, yellow) regimes. The black dashed lines show the linear fit to late-time data. The inset shows a magnified view of the initial non-linear growth. All data are for a single-site initial operator ($|\mathcal{O}|=1$) with Hamiltonian ( Eq. 1) parameters $g=1.0$ and $h=0.1$. For KC, we did a quenched average over $500$ graph realizations.
@@
Figure 9: Evolution and initial growth of Krylov complexity, $\overline{\mathcal{C}}(t)$, for a system with $L=6$. (a)-(c) The time evolution for initial operators with increasing support, $|\mathcal{O}|=1,2,$ and 33 with $\mathcal{O}=\prod\limits^{|\mathcal{O}|}_{i=1}Y_{i}$, respectively, across multiple connectance, $\tilde{M}$. Error bars indicate the standard error of the mean. (d) A magnified view of the initial growth for the same operator sizes shown in (a)-(c). All data are presented after averaging over $500$ graph realizations with Hamiltonian parameters $g=1$ and $h=0.1$ from Eq. 1.
@@
The growth of the Lanczos coefficients, $b_{n}$, directly governs the multi-stage evolution of the Krylov complexity, $\overline{\mathcal{C}}(t)$. For a single-site initial operator, $\mathcal{O}=Y_{1}$, in the chaotic regime ($\tilde{M}=0.6$), we observe that the average coefficients exhibit a power-law growth, $\overline{b_{n}}\propto n^{\delta}$, as shown in Fig. 8(a). With increasing system size $L$, the exponent $\delta$ approaches $1$, indicating the near-linear growth of $\overline{b_{n}}$ that is characteristic of chaotic systems [66, 68]. We observe an even-odd fluctuation about the linear trend in the graph averaged Lanczos coefficients ($\overline{b_{n}}$). Such fluctuations have been observed in quantum field theories [71] as well as Hamiltonian [72, 73] systems earlier. This linear growth of $b_{n}$ is known to cause an initial exponential growth of complexity for times $t\lesssim\log(L)$, a regime physically attributed to the operator scrambling across the system’s real-space support [70, 74]. For the system size studied ($L=6$), the sub-linear growth in $b_{n}$ results in an initial growth of $\mathcal{C}(t)$ that is slower than this ideal exponential behavior. This initial phase is followed by a robust linear growth, indicated by the linear fits (dashed lines) in Fig. 8(b), during which the operator spreads through the higher-order elements of the Krylov basis. Finally, at late times, the finite size of the system halts this spread, causing the complexity to saturate. These qualitative features i.e., initial non-linear growth that transitions into a linear region in KC, and saturation at late time due to finite $L$ are present in both the chaotic and near-integrable regimes, though the initial growth rate of KC slows down near integrability.
@@
The late-time saturation of Krylov complexity, $\overline{\mathcal{C}}(t\rightarrow\infty)$, also serves as a powerful indicator of chaos. The saturation value is expected to be higher for chaotic systems compared to integrable or localized systems, hinting at the emergence of more complicated operators at late times where dynamics are not hindered by constraints or approximate conserved charges. This is evident in Fig. 9(a-c), where the KC saturates to a significantly larger value (in the range of 100014001000-1400) for intermediate $\tilde{M}$, compared to the extreme $\tilde{M}$ where $\overline{\mathcal{C}}(t\rightarrow\infty)\approx 5-200$ for all operator sizes $|\mathcal{O}_{0}|$. The saturation to a value below the theoretical maximum of $K/2$ can be attributed to finite system size and potential degeneracies in the spectrum. We do not observe a clear monotonic dependence of the saturation value on $|\mathcal{O}|$, and these results do not change qualitatively for other initial operators, as shown in Appendix B.
@@
Further, as shown earlier, the initial growth rate of complexity also distinguishes the different dynamical regimes. Figure 9(d) shows that the rate of initial growth is faster in the chaotic regime compared to the near-integrable limits. The growth rate also increases with the support of the initial operator, $|\mathcal{O}_{0}|$. This is physically plausible, as an operator with a larger initial support has already “explored” a larger part of the system’s real space.
@@
Figure 10: The late-time Krylov wavefunction, $|\phi_{n}(t)|^{2}$, shown at $t=2\times 10^{5}$ for several connectance, $\tilde{M}$. The inset provides a magnified view of the occupancy for the first few Krylov basis states at the late time. All data are for a single-site initial operator ($\mathcal{O}=Y_{1}$) on a system of size $L=6$ with Hamiltonian parameters $g=1.0$ and $h=0.1$, averaged over $500$ graph realizations.

### v2

@@ IV Partial Spectral Form Factor
To complete our analysis of the pSFF, we investigate the fluctuation in pSFF due to the choice of subsystem. We compute pSFF for $L_{A}=2$ subsystems over $200$ graph realizations for multiple $L$. In each realization, we choose all possible subsystems consisting of two sites as $A$ and the rest of the sites constitute $B$. Figure 9(a) shows the distribution of asymptotic pSFF values for $L_{A}=2$ and multiple $L$ at $\tilde{M}=0.6$. The variance reduces with increasing $L$. At large $L$, the choice of sites constituting the subsytem does not affect the asymptotic pSFF in the chaotic regime. The peak also shows a shift toward the expected value (in chaotic spectra) $\approx 1/d_{A}^{2}$ ($0.0625$ for $L_{A}=2$) with increasing $L$. This behavior contrasts sharply with the near-integrable regime ($\tilde{M}=0.9$). Here, Fig. 9(b) shows a broad distribution for small $L$. The distribution gets sharper with $L$ for fixed $L_{A}$. However, it is significantly broader compared to the chaotic case for the same $L$ and $L_{A}$.
Experimental implications :
@@
Qualitative features of pSFF across dynamical regimes resemble those of SFF. As the pSFF can be experimentally obtained using randomized measurement (with lower resource requirement compared to SFF), we briefly comment on the number of measurements required to estimate pSFF with sufficient accuracy to capture the physics described in this section. As established in Ref. [52], the relative error, $\epsilon$, of the pSFF measurement scales with the number of experimental shots, $N_{m}$, as $\epsilon\propto\sigma_{A}/{K_{A}\sqrt{N_{m}}}$ at each instant of time, where $\sigma_{A}$ is the variance in the shadow estimator of pSFF. This implies that the number of measurements needed to reach a desired precision, $N_{m}\propto\sigma_{A}^{2}/K_{A}^{2}$. The number of required measurement scales is inversely proportional to the instantaneous value of pSFF. This makes the estimation of the dip in pSFF (i.e., around $t=t_{\rm Th}$) challenging for a large subsystem in a chaotic regime as $K_{A}(t\approx t_{\rm Th})=O(10^{-3})$ in our model. Compared to this, the asymptotic value of pSFF serves as a better distinguishing factor, as the magnitude is typically one order of magnitude larger here ($t>t_{p}$). However, obtaining plateau requires long time evolution, making it susceptible to decoherence. The correlation hole appears at an earlier time, though it requires a larger number of measurements.
The fluctuation of the late-time pSFF over time, within a single experimental realization, provides a qualitative diagnostic that distinguishes chaotic from integrable Hamiltonian dynamics. The markedly lower variance of the asymptotic pSFF over subsystem choices in the chaotic regime further implies a reduced experimental overhead, requiring fewer independent realizations for accurate estimation of the pSFF relative to its integrable counterpart.
@@ V Operator delocalization through Krylov complexity
In a closed quantum system, generic Hamiltonian evolution causes initial operators to become increasingly complex over time. While probes like out-of-time-ordered correlators (OTOCs) [68] can be studied for operator spreading, their reliance on spatial separation metric is ill-suited for random graphs that lack an inherent local structure. A locality-independent quantification of complexity can be the measure of the extent to which an operator spreads under time evolution in a basis composed of increasingly complicated operators [101, 95, 80]. In this section, we analyze the dynamics of operator spreading in the system through Krylov complexity. The time evolution of an operator is generated by applying the propagator constructed with the Liouvillian superoperator, and the resulting KC dynamics provide a diagnostic for quantifying chaos and localization.
@@
where the Liouvillian, $\mathcal{L}=(\mathbb{I}\otimes H-H^{T}\otimes\mathbb{I})$ and $|\mathcal{O})=\mathrm{vec}(\mathcal{O})\in\mathcal{H}_{d}\otimes\mathcal{H}_{d}$ refers to the vectorized operator defined using,
@@
The dimension of $\mathcal{K}_{\mathcal{O}}$ is denoted by $K$ and bounded [82] as $1\leq K\leq d^{2}-d+1$. Evidently, the nested commutator operators in $\mathcal{K}_{\mathcal{O}}$ gets wider (have larger support) with increasing order. We define an inner product on operator space as $(A|B)=\mathrm{Tr}(A^{\dagger}B)$, and work with a complete orthonormal basis $\{|\mathcal{O}_{n})\}_{n=0}^{K-1}$ in $\mathcal{K}_{O}$ obtained using Lanczos algorithm and implemented through successive full Gram-Schmidt orthogonalization (see Appendix D) starting from $|\mathcal{O}_{0})=\mathrm{vec}(\mathcal{O}(0))$ i.e., the first element of the basis is the initial operator. This orthonormal basis in operator space is denoted as Krylov basis. The Liouvillian takes a tri-diagonal form when expanded in $\{|\mathcal{O}_{n})\}$,
@@
where diagonal elements are 0 and off-diagonal elements ($b_{1},b_{2},\ldots b_{n}$) form an ordered set of coefficients referred to as Lanczos coefficients. Under time-evolution with $\mathcal{L}$,
@@
where $n\in[0,K-1]$ and $\phi_{n}(0)=\delta_{n,0}$ is taken as initial condition, $b_{0}=b_{K}=0$ is the boundary condition. The form of $\mathcal{L}$ suggests an interpretation of this dynamics analogous to a particle hopping problem on a finite chain, with $\{|O_{n})\}_{n=0}^{K-1}$ acting as the sites and the hopping amplitudes given by the Lanczos coefficients. To elucidate this, we can represent $\mathcal{L}$ in Krylov basis,
$\mathcal{L}=\sum_{n=0}^{K-1}b_{n}(|\mathcal{O}_{n})(\mathcal{O}_{n-1}|+|\mathcal{O}_{n-1})(\mathcal{O}_{n}|).$ (22)
@@
Figure 10: Analysis of the Lanczos coefficients, $b_{n}$, for an initial operator with support on a single site. The main panel shows the average coefficients, $\overline{b_{n}}$, as a function of the iteration number $n$ for $\tilde{M}=0.6,0.9$. Error bars represent the standard error of the mean. ( Inset )(a) shows the initial growth of $\overline{b_{n}}$ for the first 20 iterations. (b) shows the distribution of the logarithm of the ratio of consecutive coefficients, $\log(\overline{x})$, where $x_{n}=b_{n}/b_{n+1}$. (c) shows the variance of the $\log(\overline{x})$ distribution as a function of $\tilde{M}$. All data are for a system of size $L=6$ with Hamiltonian parameters $g=1.0$ and $h=0.1$, averaged over $500$ graph realizations (denoted by overbar).
@@
The corresponding evolution of wavefunction representing this particle, $\mathbb{\phi}=(\phi_{0},\imath\phi_{1},\ldots,\imath^{K-1}\phi_{K-1})$ can be obtained by solving the set of coupled equations in 21. The KC is defined as,
$\mathcal{C}(t)=\sum_{n=0}^{K-1}n|\phi_{n}(t)|^{2},$ (23)
@@
measuring the particle’s mean position in the Krylov chain. $|\mathcal{O}_{n})$ involves $n$-th order nested commutators, resulting in Krylov basis elements that have increasingly larger support and more complicated structure with increasing $n$. KC captures the spread of a time-evolved operator in this Krylov basis. We denote the support of an operator $\mathcal{O}$ as $|\mathcal{O}|$.
@@
The Lanczos coefficients $b_{n}$ control how quickly operators spread through the Krylov basis during time evolution. The scaling of $b_{n}$ with $n$ and its fluctuation serve as a signature that can distinguish quantum chaotic dynamics from integrable or localized dynamics [100]. In the following discussion, we take the initial operator in simple product form,
$\mathcal{O}(0)=\prod_{i=1}^{|\mathcal{O}|}Y_{i}.$ (24)
Here, $|\mathcal{O}|$ denotes the size of $\mathcal{O}$, i.e., the number of sites where the operator acts non-trivially. Note that we do not have any underlying spatial structure in the lattice; thus, no natural local structure can be attributed to $\mathcal{O}(0)$ when it has support larger than $1$. In the rest of the section, we present the empirical observations regarding the Krylov complexity and Lanczos coefficients as well as comparisons with related similar studies and other systems.
Fluctuation in Lanczos coefficients :
The sequence of Lanczos coefficients $b_{n}$ has been investigated across different dynamical regimes [11, 77]. Empirically, it is found that $b_{n}$} grows linearly (sub-linearly) at small $n$ for chaotic (integrable) systems, saturates at intermediate $n$, and finally descends till n=K1n=K-1. The inset (a) in Fig. 10 shows the growth and subsequent saturation of Lanczos coefficients for small $n$. The distinction in the scaling of Lanczos coefficients at small $n$ between chaotic and integrable regimes is unclear here due to the small system size ($L=6$).
The descent region in $b_{n}$ shows significantly larger fluctuations in the integrable regime than in the chaotic regime. In a single-particle picture, this translates into greater disorder in local hopping parameters. In Fig. 10 (main panel), we show the Lanczos sequence, $\overline{b_{n}}$ (averaged over ER graphs) for $\tilde{M}=0.6$ and $0.9$. The fluctuation in Lanczos amplitude at large $n$ is significantly larger for $\tilde{M}=0.9$ (approaching integrability) compared to $\tilde{M}=0.6$ (chaotic). Similar large fluctuations were also observed in the sparsely connected regime (not shown). To quantify this fluctuation, we use the distribution of the ratio of consecutive Lanczos coefficients, $x_{n}=\left(\frac{b_{n}}{b_{n+1}}\right)$ averaged over graph realizations. Figure 10 (b) shows the distribution of $\log(\bar{x}_{n})$ with $n$ for multiple $\tilde{M}$. The distribution has a mean 0 with the variance increasing as $\tilde{M}\to 1$ (Fig. 10 (c)).
@@
Figure 11: First few Lanczos coefficients and Early-time growth of Krylov complexity. (a) Log-log plot of the average Lanczos coefficients, $\overline{b_{n}}$, versus the iteration number $n$ for several system sizes $L$ at $\tilde{M}=0.6$. Dashed lines show power-law fits, $\overline{b_{n}}\propto n^{\delta}$. All data points are averaged over $200$ graph realizations. (b) The corresponding early-time growth of Krylov complexity, $\overline{\mathcal{C}}(t)$, for the chaotic ($\tilde{M}=0.4$, red) and near-integrable ($\tilde{M}=0.9$, yellow) regimes. The black dashed lines show the linear fit to late-time data. The inset shows a magnified view of the initial non-linear growth. All data are for a single-site initial operator ($|\mathcal{O}|=1$) with Hamiltonian ( Eq. 1) parameters $g=1.0$ and $h=0.1$. For KC, we performed a quenched average over $500$ graph realizations (denoted by overbar).
@@
Figure 12: Evolution and initial growth of Krylov complexity, $\overline{\mathcal{C}}(t)$, for a system with $L=6$. (a)-(b) The time evolution for initial operators with increasing support, $|\mathcal{O}|=1$ and $2$ with $\mathcal{O}=\prod\limits^{|\mathcal{O}|}_{i=1}Y_{i}$, respectively, across multiple connectance, $\tilde{M}$. Error bars indicate the standard error of the mean. (c) The asymptotic $\overline{\mathcal{C}}$ as a function of $\tilde{M}$ for $|\mathcal{O}|=1,2$ and 33 (marked with solid,dashed and dotted black lines, respectively). (d) A magnified view of the initial growth for the same $|\mathcal{O}|=1,2$ and $3$ for $\tilde{M}=0.6$ and 0.9. All data are presented after averaging over $500$ graph realizations (denoted by overbar) with Hamiltonian parameters $g=1$ and $h=0.1$ from Eq. 1.
Growth of KC :
The growth of the Lanczos coefficients, $b_{n}$, directly governs the multi-stage evolution of the Krylov complexity, $\overline{\mathcal{C}}(t)$. Figure 11(a) shows the $n$ dependence of $\overline{b_{n}}$ for a single-site initial operator, $\mathcal{O}=Y_{1}$ in the chaotic regime ($\tilde{M}=0.6$). $\overline{b}_{n}$ is expected to grow linearly with $n$ in the chaotic regime according to the universal operator growth hypothesis [80, 82]. Upon fitting $b_{n}$ calculated in finite systems to $n^{\delta}$, we find that $\delta$ approaches $1$ with increasing system size $L$, consistent with chaotic behavior. For times $t\lesssim\log(L)$, this linear growth of $b_{n}$ results in an exponentially growing KC, a feature associated with the spreading of operator support across the system [83, 96]. For $L=6$, we find the initial growth of $\mathcal{C}(t)$ to be slower than this expected exponential behavior (Fig. 11(b)), likely related to sublinear growth of $b_{n}$ with $n$ in smaller systems (Fig. 11(a)).
@@
This initial phase is followed by a robust linear growth, indicated by the linear fits (dashed lines) in Fig. 11(b), during which the operator spreads through the higher-order elements of the Krylov basis. Finally, at late-times, the finite-size of the system halts this spread, causing the complexity to saturate. These qualitative features i.e., initial non-linear growth that transitions into a linear region in KC, and saturation at late-time due to finite $L$ are present in both the chaotic and near-integrable regimes, though the initial growth rate of KC slows down near integrability (Fig. 11(b) inset).
We observe an even-odd fluctuation about the linear trend in $\overline{b_{n}}$ similar to what has been observed in quantum field theories and Hamiltonian systems [4, 84, 99, 10]. The precise form of the subleading corrections to the linear growth relates the long-time decay of the operator autocorrelation. Corrections with an odd-even structure of the form $(-1)^{n}(\ln n)^{-a}$ have been shown to imply a power-law decay of the autocorrelation [80, 10]. In the finite data in Fig. 11(a), it is unclear if the corrections scale this way.
Saturation value of KC:
@@
Among the features in the time evolution of KC, late-time saturation value $\overline{\mathcal{C}}(t\rightarrow\infty)$, is the clearest indicator of chaos. The saturation value is expected to be higher for chaotic systems compared to integrable or localized systems, hinting at the emergence of more complex operators at late-times where dynamics are not hindered by constraints or approximate conserved charges. This is evident in Fig. 12(a)-(b), where the KC saturates to a significantly larger value (in the range of 100014001000-1400) for intermediate $\tilde{M}$, compared to the extreme $\tilde{M}$ where $\overline{\mathcal{C}}(t\rightarrow\infty)\approx 5-200$ for all operator sizes $|\mathcal{O}_{0}|$. Figure  12(c) shows the KC value at saturation as a function of $\tilde{M}$, clearly indicating the distinct nature of operator spreading in the chaotic and integrable/localized regimes. The saturation to a value below the theoretical maximum of $K/2$ can be attributed to finite system size and potential degeneracies in the spectrum.
Effect of initial operator size:
@@
Figure 12(d) shows that the initial growth rate increases with the support of the initial operator, $|\mathcal{O}_{0}|$. This is physically plausible, as an operator with a larger initial support has already “explored” a larger part of the system’s real space. This feature holds in both chaotic and near-integrable regimes.
We do not observe a simple monotonic dependence of the KC saturation value on the initial operator support $|\mathcal{O}|$, and these results do not change qualitatively for other initial operators, as shown in Appendix E. In contrast, in studies on $\mathrm{SYK}_{2}$ (a free model), the growth of $\mathcal{C}(t)$ was found to depend on initial operator size [55].
@@
Figure 13: The late-time Krylov wavefunction, $|\phi_{n}(t)|^{2}$, shown at $t=2\times 10^{5}$ for several connectance, $\tilde{M}$. (Inset) A magnified view of the occupancy for the first few Krylov basis states at the late-time. All data are for a single-site initial operator ($\mathcal{O}=Y_{1}$) on a system of size $L=6$ with Hamiltonian parameters $g=1.0$ and $h=0.1$, averaged over $500$ graph realizations (denoted by overbar).
Wavefunction spread in Krylov space: